OpenStudy (vishweshshrimali5):

Let S be the sum of all two digit numbers each of which contains one odd and one even digits. Then S is divisible by (a) 2 (b) 3 (c) 5 (d) 7 There may be more than 1 correct answer

OpenStudy (vishweshshrimali5):

myninaya can u help me ?

myininaya (myininaya):

S=x+y x=10n+c y=10d+k since x and y are two digit numbers n,d,c, and k are the digits of x and y n is the 10's digit for x c is the 1's digit for x d is the 10's digit for y k is the 1's digit for y ok still thinking...

OpenStudy (vishweshshrimali5):

ok let me read this

myininaya (myininaya):

ok so n or c will be even but not both and d or k will be even but not both

myininaya (myininaya):

lets just assume for right n is even so n=2t so we have x=10(2t)+c=20t+c so c is odd so c=2s+1 x=20t+(2s+1)=20t+2s+1=2(10t+s)+1 so x is odd

OpenStudy (vishweshshrimali5):

ok

myininaya (myininaya):

ok lets assume d is even so d=2j then k=2g+1 so y=10(2k)+2g+1=2(10k+g)+1 so y is odd so S=x+y=2(10t+s)+1+2(10k+g)+1=2(10t+s)+2+2(10k+g)=2(10t+s+1+10k+g) so S is even and therefore 2 goes into S

OpenStudy (vishweshshrimali5):

I have the answer of this question but can u explain it

myininaya (myininaya):

ok

OpenStudy (anonymous):

it is certainly divisible by 2

myininaya (myininaya):

yes which i showed above

myininaya (myininaya):

with alot of different variables lol

OpenStudy (anonymous):

heh, you needn't have done all that to prove it, but it is formal math.

OpenStudy (anonymous):

wait, is 0 considered even?

myininaya (myininaya):

0 is even

myininaya (myininaya):

i define anything divisble by 2 to be even

OpenStudy (anonymous):

in the number 10, 1 is odd, but is 0 even? and are we allowed to consider the numbers 10,20,30 and so on?

OpenStudy (anonymous):

sorry, not 20

OpenStudy (vishweshshrimali5):

I have the answer of this question but can u explain it

OpenStudy (vishweshshrimali5):

The sum of all two digit numbers = 10 + 11 + 12 + .... + 99 = 99(10 + 99)/2 = 4905 There are 5 x 5 = 25 numbers with two digits. The sum of these numbers = 5(1+3+5+7+9) x 11 = 55 x 25 = 1375. Likewise, th sum of numbers with two even digits = 5(0+2+4+6+8) x 11 - (0+2+4+6+8) = 54 x 20 = 1080 Therefore, S = 4905 - 1375 = 2450 = \[2.5^2.7^2\] Therefore, correct answers are (a),(c) and (d).

OpenStudy (vishweshshrimali5):

I have the answer of this question but can u explain it

OpenStudy (vishweshshrimali5):

Now will you please explain it.

myininaya (myininaya):

wow never seen a question like i dont believe or a solution like that i mean but let me look at it

myininaya (myininaya):

dhat do you understand it

OpenStudy (vishweshshrimali5):

I have the answer of this question but can u explain it

OpenStudy (vishweshshrimali5):

what do you mean to say myininaya ? I can't understand what u mean to say.

OpenStudy (anonymous):

no, not really. I am trying to read it carefully

OpenStudy (vishweshshrimali5):

I have the answer of this question but can u explain it

OpenStudy (vishweshshrimali5):

ok

OpenStudy (vishweshshrimali5):

I have the answer of this question but can u explain it

OpenStudy (vishweshshrimali5):

Take ur time

OpenStudy (anonymous):

I understnd that there are 25 numbers with both digits odd and 20 numbers with both even digits. but I don't understand how they came up with their sum

OpenStudy (vishweshshrimali5):

I have the answer of this question but can u explain it

OpenStudy (vishweshshrimali5):

do u mean this step = 90/2 ( 10+ 99) ? I know how it came

OpenStudy (vishweshshrimali5):

I have the answer of this question but can u explain it

OpenStudy (vishweshshrimali5):

by applying the formula of sum of arithmetic progression

OpenStudy (anonymous):

no not that. i meant how they came up with 1375 and 1080

OpenStudy (vishweshshrimali5):

I have the answer of this question but can u explain it

OpenStudy (vishweshshrimali5):

sorry that i also don't know

OpenStudy (anonymous):

11+13+15+17+19+31+33+35+37+39+51+53+55+57+59+71+73+75+77+79+91+93+95+97+99

OpenStudy (anonymous):

if you add all that up you get 1375. but, I don't understand the formula you gave for that. there must be a pattern we are missing

OpenStudy (vishweshshrimali5):

I have the answer of this question but can u explain it

OpenStudy (vishweshshrimali5):

u mean 5(1 + 3 + 5 + 7 +9 ) * 11

OpenStudy (anonymous):

yes

OpenStudy (vishweshshrimali5):

I have the answer of this question but can u explain it

OpenStudy (vishweshshrimali5):

i don't know sry

OpenStudy (vishweshshrimali5):

I have the answer of this question but can u explain it

OpenStudy (vishweshshrimali5):

bhaiya batao naa please'

OpenStudy (anonymous):

well, empirically, the answer is correct. you can add up all the numbers I provided and end up with 1375 similarly if you add up 20+22+24+26+28+40+42+44+46+48+60+62+64+66+68+80+82+84+86+88 you get 1080. use your calculator. but this is not an elegant method. you are using brute force.

myininaya (myininaya):

maybe amistre can help i will give you a medal amistre :)

OpenStudy (amistre64):

Let S be the sum of all two digit numbers each of which contains one odd and one even digits. is that "at least one odd or one even"? or only one odd and one even?

OpenStudy (vishweshshrimali5):

I have the answer of this question but can u explain it

OpenStudy (vishweshshrimali5):

amisre will u ?

OpenStudy (vishweshshrimali5):

I have the answer of this question but can u explain it

OpenStudy (vishweshshrimali5):

no only 1 odd and 1 even

myininaya (myininaya):

i dont like that solution and i cant explain it :(

OpenStudy (anonymous):

anyway, if the sum of all numbers between 10 and 99 is 4905 and the sum of all numbers with both digits odd is 1375 and the sum of all numbers with both digits even is 1080, then the sum of all numbers with one odd digit and one even digit has to be 4905-1080-1375 = 2450 which is divisible by 2,5 and 7.

OpenStudy (amistre64):

1,3,5,7,9; 2,4,6,8 is 0 even or odd or neither?

OpenStudy (vishweshshrimali5):

I have the answer of this question but can u explain it

OpenStudy (vishweshshrimali5):

in my view it is even

myininaya (myininaya):

0 is even

OpenStudy (amistre64):

5*5 = 25 number that are odd then even 4*5 = 20 numbers that are even then odd right?

OpenStudy (vishweshshrimali5):

I have the answer of this question but can u explain it

OpenStudy (vishweshshrimali5):

right but how do u get that ?

myininaya (myininaya):

oh i misread the question lol i guess that could be my problem

OpenStudy (amistre64):

that many you start with how many options you get to start with; then each option has the same amount of branches from it that equal the second set of options 5 odd branches with 5 even fingers each = 25 choices

OpenStudy (vishweshshrimali5):

I have the answer of this question but can u explain it

OpenStudy (vishweshshrimali5):

k

OpenStudy (amistre64):

4 even branches with 5 odd fingers = 20 choices 45 numbers total in the pool

OpenStudy (vishweshshrimali5):

I have the answer of this question but can u explain it

OpenStudy (vishweshshrimali5):

yaaaaa

OpenStudy (amistre64):

12 14 --- 26 16 --- 42 18 --- 60 12, 26, 42, 60, 80, B{n} = B{n-1} +B{n-2} B{1} = 12 B{2} = 12 + 12 +2 B{3} = 12 + 14 + 12 +4 B{4} = 12 + 14 + 16 +12+6 B{n} = B{n-1} + B{0}+2(n-1) might be helpful

OpenStudy (amistre64):

B{1} not B{0} :)

OpenStudy (amistre64):

B{n} = B{n-1} +B{1} +2(n-1) B{n-1} = B{n-2} +B{1} +2(n-2) B{n-2} = B{n-3} +B{1} +2(n-3) B{n} = [[B{n-3} +B{1} +2(n-3)] +B{1} +2(n-2)] +B{1} +2(n-1) = B{n-3} +3B{1} +2(3n -3 -2 -1) = B{n-r} +rB{1} +2(rn -(r)-(r-1)-(r-2)-...-3-2-1) n-r = 1 when r = n-1 B{n} = B{n-(n-1)} +(n-1)B{1} +2((n-1)n -(n-1)-(n-1-1)-(n-1-2)-...-3-2-1) = 12 +(n-1)(12) +2(n^2-n -n+1-n+2-n+3-...-3-2-1) = 12(n) +2(n^2-n + A) A=(-n+1)+(-n+2)+(-n+3)+...+ (-3) + (-2) + (-1) ) A= (-1) + (-2) + (-3) +...+(-n+3)+(-n+2)+(-n+1)) ------------------------------------------------- 2A= -n + -n + -n +...+ -n + -n + -n A = -n(....)/2

myininaya (myininaya):

1 | 0 2 4 6 8 2 | 1 3 5 7 9 3 | 0 2 4 6 8 4 | 1 3 5 7 9 5 | 0 2 4 6 8 6 | 1 3 5 7 9 7 | 0 2 4 6 8 8 | 1 3 5 7 9 9 | 0 2 4 6 8 Let M be the sum of all two digit numbers M=10+11+12+....+99=(1+2+3+....+9)+10+11+12+13+...+99-(1+2+3+...+9) =99(99+1)/2-45=4950-45=4905 Look at the chart above . the only thing that is missing are: 11,13,15,17,19,20,22,24,26,28,31,33,35,37,39,40,42,44,46,48,51,53,55,57,59,60,62, 64,66,68,71,73,75,77,79,80,82,84,86,88,91,93,95,97,99 now we can separate these numbers into two groups both even digits:={20,22,24,26,28,40,42,44,46,48,60,62,64,68,80,82,84,86,88} both odd digits:={11,13,15,17,19,31,33,35,37,39,51,53,55,57,59,71,73,75,77,79,91,93,95,97,99} the sum of the elements of the both even digit set is =20+22+24+26+28+40+42+44+46+48+60+62+64+68+80+82+84+86+88 =1080 the sum of the elements of the both odd digit set is =11+13+15+17+19+31+33+35+37+39+51+53+55+57+59+71+73+75+77+79+ 91+93+95+97+99=1375 so S can be found by doing S=M-sum of both even digits-sum of both odd digits =4905-1080-1375=2450 we should write 2450 in prime factorization to better see its factors 2450=2*5^2*7^2

OpenStudy (amistre64):

B{n} = 12(n) +2n^2 -n(n+1) 12 = 12(1) + 2 - 1(2) 26 = 12(2) +2(4) - 2(3) = 24 + 8 - 6 42 = 12(3) +2(9) -3(4) = 36 + 18 -12 = 36 + 6 60 = 12(4) + 2(16) -4(5) = 48 + 32 - 20 = 48 + 12 think I got it :) at least part of it lol

myininaya (myininaya):

:)

OpenStudy (amistre64):

70 + 170 + 270 + 370 + 470 = 1350 for the odds ... maybe :)

OpenStudy (amistre64):

had to start at 10; 30; 50; 70; and 90

OpenStudy (amistre64):

2 2 2 21 41 61 23 43 63 25 45 65 27 47 67 29 49 69 ----- ---- ---- 125 225 325 425 = 1000 + 1350 = 2350 maybe?

myininaya (myininaya):

omg almost

OpenStudy (amistre64):

2 2 10 30 12 32 14 34 16 36 18 38 --- ---- 70 170 270 370 470 = 1350 right :)

OpenStudy (amistre64):

1100 + 1350 = 2450 :)

myininaya (myininaya):

yes!:)

OpenStudy (anonymous):

great amistre and myininaya u made a good effort to solve my brother's question

OpenStudy (anonymous):

thanks from my elder brother's side he would reply later

OpenStudy (anonymous):

and also dhatraditya thanks